# HIP-LLM: A Hierarchical Imprecise Probability Approach to Reliability Assessment of Large Language Models

### v1

@@ 4.1 Research Questions
RQ3 (Sensitivity to OPs): How sensitive are the posterior reliability estimates of HIP-LLM to variations in the OPs that characterize operational usage of LLMs at various levels? The delivered and perceived reliability of an LLM depends on how it will be used, i.e. the OP. General-purpose LLMs may exhibit lower sensitivity to variations in OPs, whereas LLMs trained for specific (sub-)domains are likely to be more sensitive. We investigate and demonstrate how HIP-LLM can characterize and quantify such OP-dependent reliability variations.
@@ 4.2 Experimental Setup
Since this work represents an initial academic study on reliability assessment modeling for LLMs without actual operational field data, we follow a similar approach to [31, 32] by leveraging existing benchmark datasets. To emphasize the role of OPs, we “simulate” the OPs by assigning probability distributions over these datasets for sampling and by specifying operational weights across (sub-)domains. Without loss of generality, and consistent with our three-level hierarchy, we define the task-level OP as a uniform distribution and assign operational weights at the subdomain and domain levels proportionally to their dataset sizes. That said, assessors may use alternative distributions as the 3-level OPs when additional information is available, such as those approximated from historical usage data [24, 76] or user behavior reports [27].
@@
All experiments were conducted using publicly available APIs of GPT-4o, GPT-4o-mini (from OpenAI), Claude Sonnet 4.5, and Claude Haiku 3.5 (from Anthropic) [^17] For academic research purposes only which is permitted under both vendors’ terms of service for research publications.. Table 2 reports sample accuracies per subdomain, i.e., the proportions of correct responses under Pass@1181818Produce 1 solution per task. Score as correct if that solution passes verification.. To operationalize the i.i.d. Bernoulli trail assumption on tasks, all LLM evaluations were conducted with cleared context between tasks. Each task in MBPP, DS-1000, BoolQ, and RACE-H was processed as an independent API call with no conversational history, ensuring that outcomes are not influenced by previous interactions.
@@
Dom Subdom (Data.)
4o-mini
4o
sonnet-4.5
haiku-3.5
Mean
@@ 4.3.1 RQ1 (Effectiveness)
(a) Subdomain11
(b) Subdomain12
(c) Subdomain21
(d) Subdomain22
Figure 3: Posterior CDF envelopes at subdomain level. Rows correspond to domains (D1 (coding): MBPP, DS-1000; D2 (reasoning): BoolQ, RACE-H).
@@
(a) Domain1
(b) Domain2
Figure 4: Posterior CDF envelopes for domain-level reliabilities $p_{i}=\sum_{j}\Omega_{ij}\theta_{ij}$. Operational profile weights are $\Omega_{1\cdot}=(0.204,0.796)$ for D1 and $\Omega_{2\cdot}=(0.483,0.517)$ for D2.
@@
Figure 5: Posterior CDF envelope for overall LLM reliability $p_{L}=\sum_{i}W_{i}p_{i}$ with domain weights W=(0.149, 0.851)W=(0.149,\,0.851).
@@ 4.3.2 RQ2 (Sensitivity to hyperparameters)
(a) Sensitivity of LLM posterior reliability distribution to $a_{i}$.
(b) Sensitivity of LLM posterior reliability distribution to $b_{i}$.
(c) Sensitivity of LLM posterior reliability distribution to $c_{i}$.
(d) Sensitivity of LLM posterior reliability distribution to $d_{i}$.
Figure 6: Effect of variations in the hyperparameter values ($a$, $b$, $c$, and $d$) on the overall LLM reliability.
[…]
Fig. 6(a) illustrates the sensitivity of the LLM-level posterior reliability distributions to the hyperparameter $a_{i}$, which serves as the first shape parameter of the Beta prior $\mu_{i}\sim\mathrm{Beta}(a_{i},b_{i})$. Increasing $a_{i}$ effectively shifts the prior mean of $\mu_{i}=a_{i}/(a_{i}+b_{i})$ upward, representing stronger prior belief in higher domain reliability. The rightward shift indicates a higher central estimate of LLM-level reliability.
Fig. 6(b) presents the corresponding sensitivity analysis for the hyperparameter $b_{i}$, the second shape parameter of the same Beta prior on $\mu_{i}$. Increasing $b_{i}$ decreases the prior mean $\mu_{i}=a_{i}/(a_{i}+b_{i})$, thereby encoding a more pessimistic prior expectation of domain reliability. The leftward displacement indicates a reduction in the posterior mean reliability.
Fig. 6(c) and Fig. 6(d) together illustrate how the LLM-level posterior reliability distribution responds to changes in the hyperparameters $c_{i}$ and $d_{i}$, which jointly define the Gamma prior $\nu_{i}\sim\mathrm{Gamma}(c_{i},\text{rate}=d_{i})$ controlling the pooling strength among subdomains within a domain. Increasing $c_{i}$ raises the expected prior strength $E[\nu_{i}]=c_{i}/d_{i}$ and reduces its relative variance, enforcing stronger coupling across subdomains so that their reliabilities $\theta_{ij}$ concentrate more closely around the domain mean $\mu_{i}$; this yields a modest rightward shift in the cumulative distribution, indicating a more stable and slightly higher central reliability estimate. Conversely, increasing $d_{i}$ lowers $E[\nu_{i}]$ and weakens pooling, permitting greater variability among subdomains and consequently broadening and shifting the posterior envelopes leftward. This broader, left-shifted distribution reflects increased epistemic uncertainty and the greater influence of weaker subdomains on the aggregate, leading to more conservative and less certain reliability assessments.
@@ 4.3.3 RQ3 (Sensitivity to OPs)
(a) Anthropic Sonnet 4.5
(b) OpenAI 4-o
Figure 7: Posterior CDF envelopes for Reasoning-domain ($D_{2}$) reliability under alternative operational weights on RACE–H, $\Omega_{\mathrm{RACE}}\in\{0.10,\,0.517,\,0.90\}$ with $\Omega_{\mathrm{BoolQ}}=1-\Omega_{\mathrm{RACE}}$. (a) Sonnet 4.5; (b) OpenAI 4-o.
Fig. 7 shows posterior CDF envelopes for the Reasoning-domain reliability $p_{2}$ under three operational profiles, $\Omega_{\mathrm{RACE}}\in\{0.10,\,0.517,\,0.90\}$ with $\Omega_{\mathrm{BoolQ}}=1-\Omega_{\mathrm{RACE}}$. Fig. 7(a) reports Anthropic Sonnet 4.5; Fig. 7(b) reports OpenAI 4-o.
(a) Sonnet 4.5. As $\Omega_{\mathrm{RACE}}$ increases the CDF shifts slightly left (lower delivered $p_{2}$), with strong overlap among bands. This weak sensitivity to the operational profile is consistent with the subdomain accuracies being both high and similar ($\theta_{\text{BoolQ}}\approx 0.90$ vs. $\theta_{\text{RACE-H}}\approx 0.84$): the mixture $p_{2}=\Omega_{\mathrm{BoolQ}}\theta_{\text{BoolQ}}+\Omega_{\mathrm{RACE}}\theta_{\text{RACE-H}}$ changes only modestly as weight moves from BoolQ to RACE–H. The mild widening of the envelope at higher $\Omega_{\mathrm{RACE}}$ reflects slightly larger posterior uncertainty on RACE–H relative to BoolQ.
@@ 4.3.4 RQ4 (Predictability)
(a) Expected LLM reliability $\mathbb{E}[R_{L}(n_{F})]$ versus operational horizon $n_{F}$. Imprecise posterior envelopes show decay in reliability over consecutive operations of four models.
(b) Envelope widths (upper–lower) quantify epistemic uncertainty, peaking at intermediate horizons and shrinking as reliability saturates.
Figure 8: Reliability and uncertainty of four LLMs across operational horizons.

### v2

@@ 4.1 Research Questions
RQ3 (Sensitivity to OPs): How sensitive are the posterior reliability estimates of HIP-LLM to variations in the OPs that characterize operational usage of LLMs at various levels? The delivered and perceived reliability of an LLM depends on how it will be used, i.e. the OP. We hypothesize that general-purpose LLMs may exhibit lower sensitivity to variations in OPs, whereas LLMs trained for specific (sub-)domains are likely to be more sensitive. We investigate and demonstrate how HIP-LLM can characterize and quantify such OP-dependent reliability variations.
@@
RQ5 (Comparison to baselines): How does HIP-LLM compare to established and state-of-the-art Bayesian reliability estimators? To compare the accuracy of different models against a “ground truth” reliability, we conduct synthetic simulation experiments where the “ground truth” OP and failure probabilities of sub-domains are assumed known.
RQ6 (Failure definitions): How can HIP-LLM cope with alternative definitions of task success, such as different pass@k criteria? As per Remark 4, the definition of failures of a LLM may vary and subject to uncertainties. While we do not formally study LLM failure definitions, we demonstrate how different failure definitions can be incorporated in HIP-LLM.
RQ7 (Robustness to memory effects): How robust are HIP-LLM’s posterior estimates to violations of the i.i.d. assumption caused by memory-induced dependence during LLM evaluation? While HIP-LLM is designed for reset, single-task scenarios, real-world evaluation data may not be strictly generated under this assumption. It is therefore critical to assess whether its posterior estimates remain meaningful when independence is only approximately satisfied or partially violated by memory-induced dependencies.
RQ8 (Scalability): How does the computational time and hardware RAM cost of HIP-LLM scale with the number of domains, subdomains, hyperparameter configurations, and Monte Carlo samples? Understanding the scalability of HIP-LLM with these key parameters is essential to determine if it can be applied to large-scale evaluations in practice.
@@ 4.2 Experimental Setup
To emphasize the role of OPs, we “simulate” the OPs by assigning probability distributions over these datasets for sampling and by specifying operational weights across (sub-)domains. Without loss of generality, and consistent with our three-level hierarchy, we define the task-level OP as a uniform distribution and assign operational weights at the subdomain and domain levels proportionally to their dataset sizes. That said, assessors may use alternative distributions as the 3-level OPs when additional information is available, such as those approximated from historical usage data [22, 27] or user behavior reports [15]. As demonstrated in RQ3, where we vary the operational weights.
Remark 7 (Simulated OP from benchmarks vs. real-world OP).
Similar to many software reliability modeling studies [44], HIP-LLM assumes that the OP is specified and reflects real usage conditions. In the absence of real operational data, and following the same experimental practice as [43, 46], we leverage existing benchmark datasets to “simulate” OPs for fair comparison with baselines. Specifically, benchmark datasets are treated as sampling frames, and proportional dataset sizes are adopted as a proxy OP. This choice neither implies that OP acquisition is solved nor suggests that such proxies are suitable for real deployments; rather, it serves to demonstrate how HIP-LLM integrates OPs when they are provided and to enable comparison with existing benchmark-based methods. Estimating realistic and dynamically evolving OPs from operational data is an established problem in software reliability and is typically studied separately from reliability modeling [62, 50]. While we believe existing OP estimation techniques may be applicable to LLMs, deriving OPs from real-world LLM usage data introduces new challenges and warrants dedicated future investigation.
@@
All experiments were conducted using publicly available APIs of GPT-4o, GPT-4o-mini (from OpenAI), Claude Sonnet 4.5, and Claude Haiku 3.5 (from Anthropic) [^19] For academic research purposes only which is permitted under both vendors’ terms of service for research publications.. Table 3 reports sample accuracies per subdomain, i.e., the proportions of correct responses under Pass@1202020Produce 1 solution per task. Score as correct if that solution passes verification.. To operationalize the i.i.d. Bernoulli trail assumption on tasks, all LLM evaluations were conducted with cleared context between tasks. Each task in MBPP, DS-1000, BoolQ, and RACE-H was processed as an independent API call with no conversational history, ensuring that outcomes are not influenced by previous interactions.
@@
Dom Subdom (Data.) 4o-mini 4o sonnet-4.5 haiku-3.5 Mean
@@
Interpretation of Posterior CDF Envelopes
CDF envelopes are used to jointly represent data uncertainty and epistemic uncertainty arising from imprecise prior knowledge. Since closed-form densities are generally unavailable for weighted sums of dependent reliability parameters, posterior uncertainty is characterized numerically through CDF envelopes, which support probabilistic queries and risk-aware interpretation across all levels of the hierarchy. Unless stated otherwise, all figures in this section follow this convention.
Throughout the experimental section, where appropriate, we visualize posterior uncertainty using CDF envelopes. These CDFs are constructed from observed benchmark evaluation data, consisting of $C_{ij}$ correct responses out of $N_{ij}$ tasks for each subdomain $S_{ij}$. Depending on the level of aggregation shown, a CDF represents the posterior distribution of a non-failure probability at the subdomain level ($\theta_{ij}$), the domain level ($p_{i}$), or the overall LLM level ($p_{L}$). Domain- and LLM-level CDFs are obtained by aggregating lower-level reliability parameters using the specified OP weights.
Monte Carlo sampling is used because the hierarchical aggregation of dependent subdomain reliabilities and imprecise priors leads to posterior distributions that are analytically intractable but can be evaluated efficiently and accurately via numerical sampling.
Baseline Experimental Configuration
Unless otherwise stated, all empirical experiments are conducted using a fixed baseline configuration with $m=2$ domains, $n=2$ subdomains per domain, $K=160$ hyperparameter configurations per domain, $S=3000$ Monte Carlo samples per configuration, a $(\mu,\nu)$ integration grid of size $G=2000$, a fixed CDF evaluation grid of size $T=201$, and a capped number of LLM-level configuration pairings $K_{\text{total}}\leq 512$. Domain-level parameters $\mu_{i}$ and $\nu_{i}$ are treated as latent variables and numerically integrated over a grid, while the hyper-hyperparameters $(a_{i},b_{i},c_{i},d_{i})$ are sampled from fixed intervals, with $a_{i},b_{i}\in[1,12]$ and $c_{i},d_{i}\in[1,25]$, to represent imprecise prior knowledge.
@@ 4.3.1 RQ1(Effectiveness)
(a) Posterior CDF envelopes of the non-failure probability for $\text{Subdom}_{11}$ (MBPP dataset in $\text{Dom}_{1}$) across four models.
(b) Posterior CDF envelopes of the non-failure probability for $\text{Subdom}_{12}$ (DS-1000 dataset in $\text{Dom}_{1}$) across four models.
(c) Posterior CDF envelopes of the non-failure probability for $\text{Subdom}_{21}$ (BlooQ dataset in $\text{Dom}_{2}$) across four models.
(d) Posterior CDF envelopes of the non-failure probability for $\text{Subdom}_{22}$ (RACE-H dataset in $\text{Dom}_{2}$) across four models.
Figure 3: Posterior CDF envelopes of non-failure probability at the subdomain level. Rows correspond to domains (D1 (coding): MBPP, DS-1000; D2 (reasoning): BoolQ, RACE-H). Subdomains within the same domain are statistically dependent through shared domain-level hyperparameters $(\mu_{i},\nu_{i})$, where $\mu_{i}$ represents the average domain reliability and $\nu_{i}$ controls the strength of coupling (partial pooling) among subdomains. As a result, observations from one subdomain inform the inferred reliability of the others. Domains are assumed statistically independent. A question- oriented interpretation of figure: After observing the evaluation data, how do the four models differ in terms of their subdomain-level non-failure probability?
@@
(a) Posterior CDF envelopes of non-failure probability $p_{1}=\sum_{j}\Omega_{1j}\theta_{1j}$ for $\text{Dom}_{1}$ (Coding) across four LLMs, with subdomain-level operational weights $\Omega_{1\cdot}=(0.204,\,0.796)$.
(b) Posterior CDF envelopes of non-failure probability $p_{2}=\sum_{j}\Omega_{2j}\theta_{2j}$ for $\text{Dom}_{2}$ (reasoning) across four models, with subdomain-level operational weights (0.483, 0.517).
Figure 4: Posterior CDF envelopes of non-failure probability of domain-level ($p_{i}=\sum_{j}\Omega_{ij}\theta_{ij}$). A question-oriented interpretation of figure: After observing the evaluation data, how do the four models differ in terms of their domain level non-failure probability?
@@
Figure 5: Posterior CDF envelope of non-failure probability of overall LLM level ($p_{L}=\sum_{i}W_{i}p_{i}$) with domain weights W=(0.149, 0.851)W=(0.149,\,0.851) across four models. A question-oriented interpretation of figure: Engineering question: After observing the evaluation data, how do the four LLMs differ in terms of their LLM-level non-failure probability?
@@ 4.3.2 RQ2 (Sensitivity to hyperparameters)
(a) Sensitivity of posterior CDF envelops of non-failure probability ($p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$) to $a$.
(b) Sensitivity of posterior CDF envelops of non-failure probability of LLM ($p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$) to $b$.
(c) Sensitivity of posterior CDF envelops of non-failure probability of LLM ($p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$) to $c$.
(d) Sensitivity of posterior CDF envelops of non-failure probability of LLM ($p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$) to $d$.
Figure 6: Effect of variations in the hyperparameter values ($a$, $b$, $c$, and $d$ ) on the posterior CDF envelops of non-failure probability of LLM ($p_{L}=\sum_{i=1}^{m}W_{i}p_{i}$). A question-oriented interpretation of the figures: “What are the LLM-level posterior CDF envelopes of the non-failure probability as domain-level hyperparameters vary within their admissible ranges, after observing the testing data?”
Corresponding to Remark 6, the sensitivity analysis in Fig. 6 highlights that different hyperparameters encode qualitatively different aspects of expert elicitation and therefore require different levels of care when specified. Variations in $a_{i}$ and $b_{i}$ reflect differences in expert expectations about the average domain-level reliability and mainly affect the location of the posterior reliability envelopes. As such, specifying these parameters requires careful consideration of how optimistic or pessimistic prior beliefs are justified by available evidence. In contrast, the hyperparameters $c_{i}$ and $d_{i}$ encode expert beliefs about confidence and pooling strength across subdomains, and directly control the width of the posterior envelopes.
Fig. 6(a) illustrates the sensitivity of the LLM-level posterior reliability to the hyperparameter $a_{i}$ in the prior $\mu_{i}\sim\mathrm{Beta}(a_{i},b_{i})$. For example, keeping $b_{i}=4$, an expert belief corresponding to $a_{i}=8$ yields a prior mean $\mu_{i}=0.67$, while a more optimistic belief expressed by $a_{i}=16$ increases the mean to $\mu_{i}=0.80$. When propagated through the hierarchical model and updated with the same evaluation data, these beliefs lead to a rightward shift of the posterior CDF envelope, indicating a higher central estimate of $p_{L}$, while the envelope width remains similar due to the constraining effect of the data. Fig. 6(b) shows the complementary effect of varying $b_{i}$. With $a_{2}=12$, allowing $b_{2}\in[1,12]$ corresponds to prior means $\mu_{2}\in[0.50,0.92]$, whereas narrowing the range to $b_{2}\in[4,12]$ and $[8,12]$ restricts $\mu_{2}$ to $[0.50,0.75]$ and $[0.50,0.60]$, respectively.
Fig. 6(c) and Fig. 6(d) illustrate the sensitivity of the LLM-level posterior reliability distributions to the hyperparameters $c_{i}$ and $d_{i}$. Allowing wide ranges such as $c_{2},d_{2}\in[1,25]$ corresponds to admitting both weak and strong confidence scenarios, while progressively narrowing these ranges (e.g., $c_{2}\in[9,25]$ or $[18,25]$, and similarly for $d_{2}$) excludes low- or high-confidence assumptions. When propagated through the hierarchical model and updated with the same evaluation data, these changes leave the central location of the posterior CDF largely unchanged but systematically tighten or widen the envelope, indicating that $c_{i}$ and $d_{i}$ primarily control epistemic uncertainty rather than the expected reliability level. The stronger effect of $d_{i}$ occurs because it directly reduces the pooling strength $\nu_{i}$, while $c_{i}$ mainly affects how concentrated this belief is. As a result, changing $d_{i}$ more strongly weakens pooling across subdomains and leads to larger changes in the width of the posterior reliability envelopes.
@@ 4.3.3 RQ3 (Sensitivity to OPs)
(a) Posterior CDF envelopes of non-failure probability for the Reasoning domain ($\mathrm{D}_{2}$) under alternative operational weights on RACE–H, for Anthropic Sonnet 4.5.
(b) Posterior CDF envelopes of non-failure probability for the Reasoning domain ($\mathrm{D}_{2}$) under alternative operational weights on RACE–H, for OpenAI 4-o.
Figure 7: Posterior CDF envelopes of non-failure probability for the Reasoning domain ($\mathrm{D}_{2}$) under alternative operational weights on RACE–H, $\Omega_{\mathrm{RACE}}\in\{0.10,\,0.517,\,0.90\}$, with $\Omega_{\mathrm{BoolQ}}=1-\Omega_{\mathrm{RACE}}$. (a) Sonnet 4.5; (b) OpenAI 4-o. A question-oriented interpretation of figure: How does the inferred domain-level non-failure probability change as the operational profile shifts emphasis between RACE–H and BoolQ?
Fig. 7 shows posterior CDF envelopes for the Reasoning-domain reliability $p_{2}$ under three OPs, $\Omega_{\mathrm{RACE}}\in\{0.10,\,0.517,\,0.90\}$ with $\Omega_{\mathrm{BoolQ}}=1-\Omega_{\mathrm{RACE}}$. Fig. 7(a) reports Anthropic Sonnet 4.5; Fig. 7(b) reports OpenAI 4-o.
(a) Sonnet 4.5. As $\Omega_{\mathrm{RACE}}$ increases the CDF shifts slightly left (lower delivered $p_{2}$), with strong overlap among bands. This weak sensitivity to the OP is consistent with the subdomain accuracies being both high and similar ($\theta_{\text{BoolQ}}\approx 0.90$ vs. $\theta_{\text{RACE-H}}\approx 0.84$): the mixture $p_{2}=\Omega_{\mathrm{BoolQ}}\theta_{\text{BoolQ}}+\Omega_{\mathrm{RACE}}\theta_{\text{RACE-H}}$ changes only modestly as weight moves from BoolQ to RACE–H. The mild widening of the envelope at higher $\Omega_{\mathrm{RACE}}$ reflects slightly larger posterior uncertainty on RACE–H relative to BoolQ.
@@
From a practical point of view, these results show that changing the operational weights can meaningfully change the reliability experienced by users, especially for OpenAI 4-o. The operational weights represent how often different types of tasks occur in real use, so changing them corresponds to a change in user behavior or application context. For 4-o, different task mixes lead to clearly different reliability levels, meaning that the same model can appear reliable in one application but much less so in another. In contrast, Sonnet 4.5 shows relatively stable reliability across different task mixes, indicating more uniform performance. This highlights that understanding the expected task distribution is important when deploying LLMs, particularly for models whose performance varies strongly across subdomains.
@@ 4.3.4 RQ4 (Predictability)
(a) Imprecise posterior envelopes of the expected LLM reliability ($\mathbb{E}[R_{L}(n_{F})]$) versus the required number of failure-free future tasks $n_{F}$ for reliability claims. Imprecise posterior envelopes show decay in reliability over consecutive tasks across four models.
(b) Envelope widths (upper–lower) of expected reliability $\mathbb{E}[R_{L}(n_{F})]$. Each curve represents the spread (upper–lower bound) of expected reliability as the number of future operations increases. The bell-shaped profile indicates that uncertainty grows with operational horizon before decaying as overall reliability converges toward zero.
Figure 8: Expected reliability envelopes and their width as functions of number of failure-free future tasks required by assessors (i.e., $\mathbb{E}[R_{L}(n_{F})]$ and [upper bound of enveloplower bound of envelop][$\text${upper bound of envelop}-$\text${lower bound of envelop}]). Question-oriented interpretation of the figure: “what is the posterior expected reliability of passing next $n_{F}$ future tasks, after seeing testing data?”